Seven times a number is the same as 12 more than 3 times the nimber.
Answer by MORUPHOSUOLALE(24)


SOLUTION
LETTHE NUMBER BE = X
FIRST STATEMENT IS 7 TIMES A NUMBER=
7(X)
SECOND STATEMENT ,TWELEVE MORE THAN THREE TIMES THE NUMBER= 12+3(X}
THEREFORE,
7X=12+3X
7X-3X=12
4X=12
X=12/4
X=3
SO THE NUMBER IS DEFINITELY 3.
MORUPH OSUOLALE,
CENTRAL PIEDMONT COLLEGE,
CHARLOTTE,NC